the length of a rectangular field is 10 meters less than 9 times the width. the perimeter is 140 meters. find the length.

Answer by checkley79(3341)

L=9W-10
2L+2W=140
2(9W-10)+2W=140
18W-20+2W=140
20W=140+20
20W=160
W=160/20
W=8 ANS. FOR THE WIDTH.
L=9*8-10
L=72-10
L=62 ANS FOR THE LENGTH.
PROOF:
2*62+2*8=140
124+16=140
140=140